OpenStudy (anonymous):

Find an exact value. tan 15°

jimthompson5910 (jim_thompson5910):

Hint: \[\large \tan\left(\frac{x}{2}\right) = \frac{\sin(x)}{1+\cos(x)}\]

OpenStudy (anonymous):

x would be 15?

jimthompson5910 (jim_thompson5910):

x would be 30 x/2 = 30/2 = 15

jimthompson5910 (jim_thompson5910):

you would use the unit circle to evaluate sin(x) = sin(30) and cos(x) = cos(30)

OpenStudy (anonymous):

the answer would be 2-sqrt3?

OpenStudy (anonymous):

\[\tan 15=\tan \left( 45-30 \right)=\frac{ \tan 45-\tan 30 }{1+\tan 45\tan 30}\] plug tan 45=1,tan 30=sqrt3

OpenStudy (anonymous):

correction tan30=1/sqrt3

OpenStudy (anonymous):

i see thank you